nightshade0773:

2=2(u–2) can some on help

Florisalreadytaken:

\[ 2=2(u–2) \] \[ 2=2u-4 \] \[ 6=2u \] \[\frac{6}{2}=u\] \[ ????=u \]

nightshade0773:

i dont under stand

JamesTDG:

@florisalreadytaken would linking the equation in a graphing calculator count as a direct answer?

Florisalreadytaken:

well, first we get rid of the paranthesis, by multiplying EVERY variable in it by 2: \[ 2=\color{lightskyblue}2(u–2) \overset{u\times 2,~ and ~-2 \times 2}{======= \Rightarrow } 2=\color{lightskyblue}{2u}−\color{lightskyblue}4 \] is that part clear?

nightshade0773:

yea

Florisalreadytaken:

great. so what we want to do is slowly solve for \(u\) step by step. we can see that we can move the \(-4\) to the other side, always with the opposite operator (opposite of negative is positive.) thats why we get: \[ 2=2u−4 ~~\Rightarrow ~~ \color{tomato}{2+4}=2u ~~ \Rightarrow ~~ \color{tomato}6=2u \] clear?

nightshade0773:

kinda the answer choice it gave me is u=3 and u=5

JamesTDG:

so, you have 2u=6, so what is multiplied by 2 to make 6?

nightshade0773:

u=3 how did i not notice that

Florisalreadytaken:

@florisalreadytaken wrote:
[color= inherit;][/color]\[\frac{6}{2}=u\] [color= inherit;][/color]\[ ????=u \]
it was in front of you haha

nightshade0773:

@florisalreadytaken wrote:
@florisalreadytaken wrote:
[color= inherit;][/color]\[\frac{6}{2}=u\] [color= inherit;][/color]\[ ????=u \]
it was in front of you haha
mean -^-